OpenStudy (anonymous):

Choose the graph which matches the following system of equations. –3x + y = 2 –2x – 2y = –5

ganeshie8 (ganeshie8):

check for x/y intercepts in the graphs

OpenStudy (anonymous):

how so?

OpenStudy (anonymous):

A graph shows two lines intersecting in the first quadrant. One line passes through points (0, –2) and ( 2, 4). The other line passes through points (2.5, 0) and (0, 2.5). A graph shows two lines intersecting in the fourth quadrant. One line passes through points (0, –2.5) and (–2.5, 0). The other line passes through points (–1, 1) and (0, –2). A graph shows two lines intersecting in the second quadrant. One line passes through points (0, –2) and (–2, 4). The other line passes through points (2.5, 0) and (0, 2.5). A graph shows two lines intersecting in the first quadrant. One line passes through points (0, 2.5) and (2.5, 0). The other line passes through points (0, 2) and ( –1, –1).

OpenStudy (texaschic101):

to find the y intercept, sub in 0 for x -3x + y = 2 -3(0) + y = 2 y = 2....the y intercept is (0,2) to find the x intercept, sub in 0 for y -3x + y = 2 -3x + 0 = 2 -3x = 2 x = -2/3 -- the x intercept is (-2/3,0) now the other line... -2x - 2y = -5 y intercept = 5/2 or 2.5.....(0,2.5) x intercept = 5/2 or 2.5....(2.5,0) They will cross in the 1st quadrant

OpenStudy (anonymous):

thank you

OpenStudy (texaschic101):

no problem :)